# A Network Science Approach to Granular Time Series Segmentation

### v1

@@ 7.3 Analysis of Training Sample Ratios on Model Performance
To evaluate the effect of the number of training samples on classification performance, we compared the proposed WDPVG-based model and the baseline model using the methodology described in Section 6. Figure 9 illustrates how the ratio of training to testing data impacts the model’s performance. The graph displays the average weighted F1 scores across all 59 datasets. As expected, the F1 scores for both the proposed and baseline models gradually improve as the proportion of training data increases. Also, this figure highlights the advantages of our approach. Even when using a fraction of only 10% of the available data for training, it achieves an impressive average F1 score of 0.95, significantly outperforming the baseline score of 0.73. This trend continues also for other ratios, which proves that our approach requires considerably less training data to achieve higher model performance, compared to the baseline.
@@
Additionally, we compared the number of trainable parameters of the proposed and basline model. As it can be seen in Table 2, our proposed model has $\approx$10 times less weights compared to the seq2point baseline, which means our proposed method is much less complex compared to the baseline model, and on average performs with a much higher F1 score across all the datasets.
[b] Model Number of trainable weights Proposed model $\approx$0.117M seq2point $\approx$1.174M
@@ Acknowledgements
Author contributions statement
B.B., C.F. and I.K. conceived the experiments, I.K. conducted the experiments, B.B. and I.K. analysed the results. M.M. and C.F. provided funding. I.K. and B.B. wrote the original draft. All authors reviewed the manuscript.
@@
In this section, we have summarized the detailed performance results in Tables 2-6, where each table groups the datasets by the same number of classes or distinct segments. The first column shows the name of the dataset, while the remaining columns show the selected metrics for the trained models, more specifically the proposed WDPVG, the Weighted NVG (Euclidean distance) and the baseline model seq2point. The last row in each table shows the average values of these metrics for each subset of the dataset. As we can see in Table 3, the WDPVG and the NVG perform quite similarly, with the NVG winning or drawing in 10/21 records and the WDPVG achieves 6 draws. The baseline model performs best overall, winning or drawing in 15/21 datasets, demonstrating its capability in handling simpler segmentation tasks with 2 segments. Although the additional information provided by the Reflected Perspective NVG contributes to the high F1 score values of the WDPVG, it does not offer a clear advantage over the other two models.
[b] Dataset (num. segm.) Proposed method NVG seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 ArrowHead (2) 0.997 0.997 0.997 1.000 1.000 1.000 1.000 1.000 1.000 Beef (2) 0.996 0.996 0.996 0.986 0.986 0.986 1.000 1.000 1.000 BeetleFly (2) 0.998 0.998 0.998 1.000 1.000 1.000 1.000 1.000 1.000 BirdChicken (2) 1.000 1.000 1.000 1.000 1.000 1.000 0.988 0.988 0.988 ChlorineConcentration (2) 0.682 0.682 0.682 0.838 0.838 0.838 0.917 0.910 0.906 Coffee (2) 0.980 0.980 0.980 0.980 0.980 0.980 0.994 0.994 0.994 Computers (2) 0.998 0.998 0.998 0.999 0.999 0.999 0.942 0.942 0.942 DistalPhalanxOutlineAgeGroup (2) 0.993 0.993 0.993 0.993 0.993 0.993 0.682 0.826 0.747 ECG200 (2) 0.992 0.992 0.992 0.991 0.991 0.991 1.000 1.000 1.000 ECGFiveDays (2) 1.000 1.000 1.000 0.993 0.993 0.993 1.000 1.000 1.000 GunPoint (2) 0.997 0.997 0.997 0.997 0.997 0.997 1.000 1.000 1.000 ItalyPowerDemand (2) 0.994 0.994 0.994 0.882 0.848 0.848 0.997 0.997 0.997 Lightning2 (2) 0.996 0.996 0.996 1.000 1.000 1.000 0.995 0.995 0.995 MedicalImages (2) 1.000 1.000 1.000 1.000 1.000 1.000 0.993 0.993 0.993 MoteStrain (2) 1.000 1.000 1.000 1.000 1.000 1.000 0.996 0.996 0.996 SonyAIBORobotSurface1 (2) 0.996 0.996 0.996 1.000 1.000 1.000 1.000 1.000 1.000 SonyAIBORobotSurface2 (2) 0.997 0.997 0.997 1.000 1.000 1.000 1.000 1.000 1.000 ToeSegmentation1 (2) 0.993 0.993 0.993 0.992 0.992 0.992 1.000 1.000 1.000 ToeSegmentation2 (2) 0.998 0.998 0.998 0.998 0.998 0.998 1.000 1.000 1.000 TwoLeadECG (2) 1.000 1.000 1.000 0.989 0.989 0.989 1.000 1.000 1.000 Yoga (2) 0.931 0.931 0.931 0.925 0.925 0.925 0.991 0.991 0.991 Mean 0.978 0.978 0.978 0.972 0.972 0.972 0.974 0.981 0.979 STD 0.072 0.072 0.072 0.049 0.049 0.049 0.069 0.042 0.056
Table 3: Classification results of proposed model compared to the baseline for 2 class segmentation.
Already in Tables 4 and 5, the WDPVG outperforms both NVG and the baseline by consistently achieving higher F1 scores across the majority of datasets.
[b] Dataset (num. segm.) Proposed method NVG seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 CBF (3) 1.000 1.000 1.000 0.989 0.989 0.989 0.994 0.994 0.994 CinCECGTorso (3) 0.995 0.995 0.995 0.999 0.999 0.999 0.865 0.865 0.865 DiatomSizeReduction (3) 1.000 1.000 1.000 1.000 1.000 1.000 0.930 0.919 0.919 DistalPhalanxTW (3) 0.973 0.973 0.973 0.963 0.963 0.963 0.410 0.640 0.500 Haptics (3) 0.997 0.997 0.997 0.974 0.974 0.974 0.910 0.908 0.906 InlineSkate (3) 0.997 0.997 0.997 0.997 0.997 0.997 0.844 0.826 0.824 LargeKitchenAppliances (3) 0.996 0.996 0.996 0.996 0.996 0.996 0.938 0.936 0.936 Meat (3) 0.989 0.989 0.989 0.974 0.974 0.974 0.939 0.935 0.934 OliveOil (3) 0.235 0.485 0.317 0.235 0.485 0.317 0.559 0.508 0.369 OSULeaf (3) 0.986 0.986 0.986 0.993 0.993 0.993 0.995 0.995 0.995 ProximalPhalanxOutlineAgeGroup (3) 0.990 0.990 0.990 0.984 0.984 0.984 0.896 0.889 0.891 Trace (3) 0.999 0.999 0.999 0.996 0.996 0.996 0.925 0.916 0.914 WordSynonyms (3) 0.998 0.998 0.998 0.992 0.992 0.992 0.738 0.859 0.794 Mean 0.935 0.935 0.935 0.929 0.929 0.929 0.842 0.861 0.834 STD 0.202 0.202 0.202 0.209 0.209 0.209 0.168 0.134 0.181
Table 4: Classification results of proposed model compared to the baseline for 3 class segmentation.
[b] Dataset (num. segm.) Proposed method NVG seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Adiac (4) 0.996 0.996 0.996 1.000 1.000 1.000 0.742 0.845 0.786 Car (4) 0.993 0.993 0.993 0.993 0.993 0.993 0.882 0.872 0.868 FaceFour (4) 0.996 0.996 0.996 0.994 0.994 0.994 0.886 0.883 0.877 FiftyWords (4) 0.999 0.999 0.999 0.992 0.992 0.992 0.966 0.960 0.961 InsectWingbeatSound (4) 1.000 1.000 1.000 1.000 1.000 1.000 0.995 0.995 0.995 Mallat (4) 0.995 0.995 0.995 0.992 0.992 0.992 0.910 0.912 0.910 ProximalPhalanxTW (4) 0.950 0.950 0.950 0.966 0.966 0.966 0.737 0.844 0.786 SwedishLeaf (4) 0.998 0.998 0.998 1.000 1.000 1.000 0.936 0.934 0.935 Symbols (4) 1.000 1.000 1.000 1.000 1.000 1.000 0.917 0.902 0.896 Mean 0.992 0.992 0.992 0.993 0.993 0.993 0.890 0.905 0.890 STD 0.015 0.015 0.015 0.010 0.010 0.010 0.067 0.048 0.067
Table 5: Classification results of proposed model compared to the baseline for 4 class segmentation.
In Table 6, we observe that our method’s capabilities are limited on gesture recognition datasets (Cricket X, Y and Z, UWaveGestureLibraryAll, X, Z), where the baseline method has a clear advantage.
[b] Dataset (num. segm.) Proposed method NVG seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 CricketX (5) 0.995 0.995 0.995 0.998 0.998 0.998 0.995 0.995 0.995 CricketY (5) 0.983 0.983 0.983 0.988 0.988 0.988 0.991 0.991 0.991 CricketZ (5) 0.991 0.991 0.991 0.983 0.983 0.983 0.996 0.996 0.996 Lightning7 (5) 0.998 0.998 0.998 0.997 0.997 0.997 0.827 0.829 0.825 ShapesAll (5) 0.999 0.999 0.999 0.997 0.997 0.997 0.992 0.992 0.992 SyntheticControl (5) 0.991 0.991 0.991 0.992 0.992 0.992 0.691 0.813 0.739 UWaveGestureLibraryAll (5) 0.973 0.973 0.973 0.965 0.965 0.965 0.994 0.994 0.994 UWaveGestureLibraryX (5) 0.977 0.977 0.977 0.987 0.987 0.987 0.991 0.991 0.991 UWaveGestureLibraryY (5) 0.981 0.981 0.981 0.975 0.975 0.975 0.962 0.960 0.959 UWaveGestureLibraryZ (5) 0.984 0.984 0.984 0.991 0.991 0.991 0.994 0.994 0.994 Mean 0.987 0.987 0.987 0.987 0.987 0.987 0.943 0.955 0.948 STD 0.008 0.008 0.008 0.010 0.010 0.010 0.097 0.068 0.085
Table 6: Classification results of proposed model compared to the baseline for 5 class segmentation.
@@
Table 7 shows that our proposed method is particularly suitable for datasets with six or more segments. It consistently outperforms the seq2point baseline (except for FaceAll) and demonstrates its effectiveness in capturing complex structural dependencies.
[b] Dataset (num. segm.) Proposed method NVG seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 NonInvasiveFetalECGThorax1 (6) 0.997 0.997 0.997 0.997 0.997 0.997 0.890 0.882 0.841 NonInvasiveFetalECGThorax2 (7) 0.992 0.992 0.992 0.995 0.995 0.995 0.947 0.943 0.942 FaceAll (7) 0.971 0.971 0.971 0.983 0.983 0.983 0.996 0.996 0.996 FacesUCR (7) 0.981 0.981 0.981 0.976 0.976 0.976 0.781 0.876 0.823 Fish (7) 0.974 0.974 0.974 0.972 0.972 0.972 0.911 0.907 0.904 Plane (7) 0.996 0.996 0.996 0.988 0.988 0.988 0.785 0.880 0.828 Mean 0.985 0.985 0.985 0.985 0.985 0.985 0.885 0.914 0.889 STD 0.010 0.010 0.010 0.009 0.009 0.009 0.079 0.043 0.064
@@
Table 7: Classification results of proposed model compared to the baseline for 6 and 7 class segmentation.

### v2

@@ 7.3 Analysis of Training Sample Ratios on Model Performance
To evaluate the effect of the number of training samples on classification performance, we compared the proposed WDPVG-based model and the baseline model using the methodology described in Section 6. Figure 9 illustrates how the ratio of training to testing data impacts the model’s performance. The graph displays the average weighted F1 scores across all 59 datasets. As expected, the F1 scores for both the proposed and baseline models gradually improve as the proportion of training data increases. Also, this figure highlights the advantages of our approach. Even when using a fraction of only 10% of the available data for training, it achieves an impressive average F1 score of 0.95, significantly outperforming the baseline score of 0.73. This trend continues also for other ratios, which proves that our approach requires considerably less training data to achieve higher model performance compared to the baseline.
@@
Additionally, we compared the number of trainable parameters between the proposed and baseline model. As shown in Table 3, our proposed model has $\approx$22 times fewer weights compared to the seq2point baseline (0.053M vs. 1.174M). This demonstrates that our graph-based approach achieves significantly higher parameter efficiency, delivering superior average F1 scores across all datasets with a fraction of the model complexity.
ll [b] Model Number of trainable weights WDPVG $\approx$0.053M seq2point $\approx$1.174M
@@
7.4 Ablation study results
Following the methodology detailed in Section 6.4, our Neural Architecture Search (NAS) identified a Pareto front of optimal architectures that effectively balance the trade-off between model complexity and performance. These non-dominated solutions are summarized in Table 4. Each entry represents a configuration where no other discovered solution achieves a higher F1-score without a simultaneous increase in parameter count.
To understand the structural properties that drive high performance along this Pareto front, we analyzed the marginal contribution of each architectural component across all search trials. These dependencies are visualized in Figure 10.
Figure 10: Marginal impact of architectural components—Depth ($L$), Attention Heads ($H$), and Hidden Dimension (WW)—on the mean F1-score across all search trials.
Crucially, the results reveal that architectural depth is the primary determinant of performance. As illustrated in the left panel of Figure 10, increasing network depth ($L$) from 2 to 5 layers yields the most significant performance gain. Models limited to a depth of $2$ layers ($L2$) averaged an F1-score of only $0.771$, whereas increasing the depth to $5$ layers ($L5$) pushed the performance to a near-perfect $0.994$. This structural necessity is reflected in the Pareto front (Table 4), where every top-performing configuration (Trials 89, 93, 66) utilizes a depth of at least 5 layers. This confirms that a deeper receptive field is essential for capturing the long-range temporal dependencies required for accurate segmentation.
Secondly, the comparison between Graph Attention (GAT) and Graph Convolutional (GCN) architectures quantifies the critical role of the attention mechanism. As shown in the center panel of Figure 10 and detailed in Table 5, the attention mechanism ($H$) significantly outperforms the non-attentive GCN baseline ($H0$), which achieved a mean F1-score of only $0.732$. Conversely, utilizing just 2 attention heads ($H2$) jumps the performance to $0.951$. This substantial gap of over 20 percentage points validates our hypothesis: the dynamic neighbor weighting provided by GATs is not merely an optimization, but a functional requirement for solving the dense labeling problem in this domain. Unlike GCNs, which aggregate information from all neighbors equally, GATs can effectively filter out noisy or irrelevant connections inherent in VGs.
Finally, the right panel of Figure 10 provides crucial insight into parameter efficiency regarding the hidden dimension width (W). While increasing the width to 64 units (W64) yields a peak mean F1-score of 0.829, outperforming the W32 configuration (0.804), the performance gain is marginal relative to the increase in model complexity. Beyond W64, the metric plateaus or even slightly degrades. Consequently, we selected the W32 configuration for the final architecture. The slight trade-off in raw performance is justified by the substantial reduction in parameter count, validating our preference for ”thin and deep” architectures that maximize efficiency without compromising the model’s ability to capture essential features.
Synthesizing these structural insights, we selected Trial 66 (Table 4) as our final optimized architecture. While Trial 93 represents the efficient ”knee” of the curve (0.0190.019M params, $0.984$ F1), Trial 66 leverages the optimal combination of depth ($5$ layers) and attention ($4$ heads) to achieve a peak F1-score of $0.995$. We deemed the increase in the number of trainable parameters acceptable given the precision gains. Notably, this optimized model remains approximately $22\times$ more parameter-efficient than the baseline seq2point model ($\approx 1.17$M parameters), validating the superiority of the graph-based approach.
Trial Layers Hidden Dim. Heads Learning Rate Weight Decay Params (M) Mean F1
68 2 32 2 $2.3\times 10^{-3}$ 0.0 0.001 0.732
78 2 32 4 $3.0\times 10^{-3}$ 0.0003 0.001 0.795
70 2 32 6 $1.3\times 10^{-3}$ 0.0003 0.002 0.821
50 2 64 6 $1.2\times 10^{-3}$ 0.0 0.003 0.829
27 3 32 2 $2.3\times 10^{-3}$ 0.0 0.005 0.898
36 4 32 2 $1.0\times 10^{-3}$ 0.0001 0.010 0.951
89 5 32 2 $5.7\times 10^{-4}$ 0.0001 0.014 0.960
93 6 32 2 $8.2\times 10^{-4}$ 0.0 0.019 0.984
66 5 32 4 $5.1\times 10^{-4}$ 0.0 0.053 0.995
Table 4: Pareto front solutions from the NAS, detailing the trade-off between model parameters and mean F1-score.
Model Mean Precision Mean Recall Mean F1-Score
GCN 0.741 0.740 0.732
GAT 0.995 0.995 0.995
Table 5: Performance comparison of the optimized GAT model against the GCN ablation baseline, averaged across all datasets.
@@ Acknowledgements
In this section, we summarize the detailed performance results presented in Tables 6 to 10, where each table groups datasets by the same number of classes or distinct segments. The first column lists the dataset names, while the remaining columns present the selected metrics for the proposed WDPVG method, the Euclidean distance-weighted NVG transformation combined with our proposed model, and the baseline seq2point model. The final two rows in each table report the Mean and Standard Deviation (STD) of these metrics for each dataset subset.
As shown in Table 6, the proposed method outperforms the other models for 2-segment tasks. It achieves the highest Mean F1 score of 0.991, compared to 0.972 for NVG and 0.976 for seq2point. The proposed method achieves a higher or equal F1 score than seq2point in 12 out of 21 datasets, and a higher or equal score than NVG in 15 out of 21 datasets. The proposed method demonstrates particular robustness on difficult datasets (such as ChlorineConcentration and DistalPhalanxOutlineAgeGroup), contributing to its superior overall Mean score.
In Tables 7 and 8, the proposed method continues to outperform both NVG and the seq2point baseline. For 3-class segmentation, shown in Table 7, it achieves a Mean F1 score of 0.945 (compared to 0.936 for NVG and 0.833 for seq2point) and secures a higher or equal F1 score than NVG in 11 out of 13 datasets and than seq2point in 10 out of 13 datasets. This trend remains strong for 4-class segmentation, shown in Table 8, where the proposed method achieves the highest F1 score in 8 out of 9 datasets compared to seq2point, yielding a near-perfect Mean F1 score of 0.996.
In Table 9, the proposed method maintains its strong performance, achieving the highest Mean F1 score of 0.993 (compared to 0.990 for NVG and 0.973 for seq2point). The baseline method achieves a higher score only on ’UWaveGestureLibraryAll’ and ’SyntheticControl’.
@@
Finally, Table 10, which includes datasets with 6 or 7 segments, confirms that our proposed method is particularly suitable for tasks with a higher number of segments. It outperforms the seq2point baseline in all 6 datasets. The proposed method achieves the highest Mean F1 score of 0.994, demonstrating its effectiveness in capturing complex structural dependencies.
Overall, the results show that the proposed method is the most reliable and high-performing model across all levels of segmentation complexity, from 2 to 7 classes.
[b] Dataset (num. segm.) Proposed method ClaSP + KMeans seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 ArrowHead (2) 0.99890 0.99888 0.99899 0.99280 0.99269 0.992695 0.97522 0.97374 0.97373 Beef (2) 0.99646 0.99645 0.99645 0.98687 0.98652 0.98652 0.99756 0.99754 0.99755 BeetleFly (2) 0.99806 0.99805 0.99805 0.98780 0.98750 0.98749 0.99803 0.99801 0.99802 BirdChicken (2) 1.00000 1.00000 1.00000 0.99689 0.99687 0.99688 0.99675 0.99661 0.99668 ChlorineConcentration (2) 0.90768 0.90284 0.90438 0.48558 0.69683 0.57233 0.91049 0.89684 0.88789 Coffee (2) 0.99670 0.99666 0.99667 0.99701 0.99700 0.99699 0.99826 0.99825 0.99824 Computers (2) 0.99617 0.99615 0.99616 0.25000 0.50000 0.33333 0.95503 0.95416 0.95415 DistalPhalanxOutlineAgeGroup (2) 0.99656 0.99650 0.99651 0.96587 0.96433 0.96307 0.66265 0.81401 0.73056 ECG200 (2) 0.99586 0.99583 0.99583 0.98572 0.98541 0.98532 0.99925 0.99923 0.99921 ECGFiveDays (2) 0.99582 0.99573 0.99576 0.99745 0.99744 0.99744 0.99778 0.99775 0.99773 GunPoint (2) 0.99476 0.99467 0.99469 0.97803 0.97707 0.97703 0.99374 0.99366 0.99360 ItalyPowerDemand (2) 0.99586 0.99591 0.99589 0.99265 0.99254 0.99253 1.00000 1.00000 1.00000 Lightning2 (2) 0.99587 0.99583 0.99581 0.44444 0.66666 0.53334 0.99571 0.99563 0.99560 MedicalImages (2) 0.99933 0.99931 0.99932 0.99136 0.99111 0.99114 0.98693 0.98695 0.98594 MoteStrain (2) 0.99901 0.99902 0.99900 0.99237 0.99226 0.99226 0.99696 0.99693 0.99695 SonyAIBORobotSurface1 (2) 0.99421 0.99405 0.99407 0.99083 0.99071 0.99067 0.99635 0.99629 0.99627 SonyAIBORobotSurface2 (2) 0.99906 0.99905 0.99905 0.98663 0.98632 0.98628 0.99709 0.99706 0.99706 ToeSegmentation1 (2) 0.99941 0.99939 0.99940 0.76467 0.59711 0.52136 0.99939 0.99937 0.99938 ToeSegmentation2 (2) 0.99785 0.99783 0.99784 0.75238 0.55832 0.45322 0.99891 0.99889 0.99890 TwoLeadECG (2) 0.98640 0.98565 0.98596 0.27279 0.52229 0.35839 1.00000 1.00000 1.00000 Yoga (2) 0.96976 0.96975 0.96973 0.8169 0.69450 0.67305 0.98986 0.98974 0.98962 Mean 0.99113 0.99084 0.99093 0.83948 0.86064 0.82292 0.97362 0.98003 0.97558 STD 0.02020 0.02121 0.02089 0.25303 0.17865 0.23354 0.07424 0.04475 0.06159
@@
Table 6: Classification results of the proposed model compared to the baselines for 2-class segmentation.
[b] Dataset (num. segm.) Proposed method ClaSP + KMeans seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 CBF (3) 0.97590 0.97396 0.97417 0.98976 0.98958 0.98955 0.99639 0.99633 0.99633 CinCECGTorso (3) 0.99699 0.99698 0.99697 0.52290 0.50857 0.43927 0.83061 0.82350 0.82386 DiatomSizeReduction (3) 1.00000 1.00000 1.0000 0.54821 0.53673 0.47361 0.81082 0.80455 0.80109 DistalPhalanxTW (3) 0.98452 0.99449 0.98447 0.73358 0.84147 0.78007 0.58136 0.73188 0.63292 Haptics (3) 0.98620 0.98608 0.98609 0.15631 0.39536 0.22405 0.92194 0.92068 0.92028 InlineSkate (3) 0.99862 0.99863 0.99861 0.45923 0.54076 0.44911 0.83934 0.83525 0.83491 LargeKitchenAppliances (3) 0.99725 0.99723 0.99724 0.49735 0.65587 0.54770 0.91888 0.91855 0.91854 Meat (3) 0.98969 0.98978 0.98959 0.62183 0.48303 0.46246 0.91099 0.90483 0.90367 OliveOil (3) 0.49697 0.40328 0.37505 0.66643 0.66086 0.63333 0.59613 0.51947 0.38795 OSULeaf (3) 0.99606 0.99598 0.99600 0.99480 0.99471 0.99472 0.99919 0.99917 0.99918 ProximalPhalanxOutlineAgeGroup (3) 0.99545 0.99541 0.99542 0.68465 0.74775 0.704415 0.91310 0.90337 0.90459 Trace (3) 0.99804 0.99802 0.99803 0.98042 0.97994 0.97986 0.91620 0.91104 0.90971 WordSynonyms (3) 0.99111 0.99095 0.99099 0.71489 0.84551 0.77473 0.73821 0.85936 0.79496 Mean 0.95437 0.94775 0.94482 0.65926 0.70616 0.65022 0.84409 0.85600 0.83292 STD 0.13761 0.16374 0.17135 0.23933 0.20172 0.24336 0.13428 0.12542 0.16479
Table 7: Classification results of proposed model compared to the baseline for 3-class segmentation.
[b] Dataset (num. segm.) Proposed method ClaSP + KMeans seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Adiac (4) 1.000 1.000 1.000 0.73414 0.67471 0.67175 0.90071 0.91911 0.90474 Car (4) 0.99547 0.99538 0.99539 0.69140 0.79611 0.72904 0.94417 0.94089 0.94077 FaceFour (4) 0.99862 0.99863 0.99861 0.98922 0.98917 0.98916 0.85255 0.88463 0.86533 FiftyWords (4) 0.99659 0.989647 0.99654 0.99333 0.99326 0.99326 0.99677 0.99676 0.99675 InsectWingbeatSound (4) 0.99616 0.99608 0.99610 0.97984 0.97890 0.97890 0.99360 0.993225 0.99325 Mallat (4) 0.99291 0.99286 0.9929 0.99376 0.99369 0.99367 0.76916 0.82100 0.78026 ProximalPhalanxTW (4) 0.98409 0.98375 0.98377 0.99133 0.99075 0.99088 0.68450 0.80717 0.73971 SwedishLeaf (4) 0.99810 0.99809 0.99808 0.91286 0.88534 0.87694 0.97865 0.97832 0.97835 Symbols (4) 0.99826 0.99825 0.99824 0.84345 0.66596 0.62701 0.94647 0.94475 0.94167 Mean 0.99558 0.99550 0.99551 0.90326 0.88532 0.87229 0.89629 0.92065 0.90454 STD 0.00478 0.00488 0.00487 0.12200 0.13437 0.14732 0.10846 0.07025 0.09264
Table 8: Classification results of the proposed model compared to the baselines for 4-class segmentation.
[b] Dataset (num. segm.) Proposed method ClaSP + KMeans seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 CricketX (5) 0.99624 0.99623 0.99622 0.99365 0.99353 0.99353 0.995 0.99616 0.99616 CricketY (5) 0.99246 0.99241 0.99243 0.98683 0.98666 0.98665 0.98274 0.98236 0.98234 CricketZ (5) 0.99516 0.99510 0.99512 0.99337 0.99337 0.99337 0.98550 0.98501 0.98492 Lightning7 (5) 0.99543 0.99542 0.99541 0.78671 0.87261 0.82217 0.86212 0.86106 0.85634 ShapesAll (5) 0.99844 0.99842 0.99843 0.99373 0.99359 0.99359 0.98105 0.98031 0.98037 SyntheticControl (5) 0.99382 0.99379 0.99377 0.99551 0.99546 0.99546 0.99955 0.99955 0.99955 UWaveGestureLibraryAll (5) 0.97951 0.97928 0.97933 0.996613 0.99609 0.99610 0.97348 0.99940 0.99940 UWaveGestureLibraryX (5) 0.99120 0.99113 0.99116 0.99336 0.99325 0.993226 0.98673 0.98676 0.97129 UWaveGestureLibraryY (5) 0.99653 0.99645 0.99647 0.98191 0.98084 0.98096 0.97348 0.97111 0.97129 UWaveGestureLibraryZ (5) 0.9930 0.99525 0.99527 0.72013 0.814407 0.75203 0.97661 0.97593 0.97585 Mean 0.99341 0.99335 0.99336 0.94418 0.96198 0.95071 0.97433 0.97377 0.97330 STD 0.00530 0.00536 0.00535 0.10185 0.06409 0.08792 0.04046 0.04076 0.04220
Table 9: Classification results of the proposed model compared to the baselines for 5-class segmentation.
[b] Dataset (num. segm.) Proposed method ClaSP + KMeans seq2point Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 NonInvasiveFetalECGThorax1 (6) 0.99639 0.99637 0.99638 0.99735 0.99734 0.99734 0.87522 0.86297 0.86213 NonInvasiveFetalECGThorax2 (7) 0.99107 0.99103 0.99102 0.99725 0.99723 0.99724 0.88387 0.90718 0.89291 FaceAll (7) 0.99675 0.99677 0.99676 0.71970 0.67371 0.62776 0.99111 0.99049 0.99051 FacesUCR (7) 0.99462 0.99460 0.99461 0.78491 0.79114 0.76723 0.83795 0.90262 0.86606 Fish (7) 0.98723 0.98714 0.98717 0.99242 0.99229 0.99230 0.88477 0.88668 0.88238 Plane (7) 0.99607 0.99603 0.99602 0.79975 0.83783 0.80558 0.85076 0.91197 0.87684 Mean 0.99369 0.99366 0.99366 0.88190 0.88159 0.86458 0.88728 0.91032 0.89514 STD 0.00379 0.00382 0.00381 0.13093 0.13627 0.15497 0.05424 0.04313 0.04803
Table 10: Classification results of the proposed model compared to the baselines for 6 and 7-class segmentation.